A rock climber wears of 7.5kg back pack while scaling a cliff. After 30.0mins the climber is 8.2m above the starting points.


(A) how much work does the climber do on the back pack?

(B) if the climber, weight 645N,how much does she do lifting herself and the back pack?

(C) what is the average power developed by the climber? 


1
Expert's answer
2021-02-28T07:32:12-0500

(A) By the definition of the work done, we have:


"W_{back\ pack}=mgh=7.5\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot8.2\ m=602.7\ J."

(B)

"W=W_{back\ pack}+W_{climber},""W=W_{back\ pack}+Fh,""W=602.7\ J+645\ N\cdot8.2\ m=5891.7\ J."

(C) By the definition of the power, we have:


"P_{avg}=\dfrac{W}{t}=\dfrac{5891.7\ J}{30\cdot60\ s}=3.3\ W."